### v2

Abstract
Mountain huts are buildings located at high altitude, offering a place for hikers and providing shelter. Energy supply to mountain huts remains an ongoing issue. Using renewable energies could be an appropriate solution. Jiujiu Cabins, a famous mountain hut in Shei-Pa National Park, Taiwan, has operated an off-grid solar energy storage system (ESS) with lead-acid batteries. In 2021, a serious system failure took place, leading to no electricity. After a detailed on-site survey, a reorganization and repair project was implemented, and the energy system came back to operate normally. Meanwhile, an eco-friendly lithium iron phosphate battery (LFP battery) ESS replaces part of the lead-acid battery ESS, forming a hybrid ESS, making a better and greener off-grid solar ESS. In this case report, the energy architecture, detailed descriptions, and historical status of the system are provided. An on-site survey of the failed energy system, a system improvement project, and a future plan are listed.
Index Terms:
microgrid, renewable energy, photovoltaic system, energy storage system, hybrid energy storage system, lithium-ion battery, lithium iron phosphate battery, high mountain, mountain hut, high altitude, engineering.
[…]
=-21pt
I Introduction
Mountain huts are buildings located at high altitude, in mountainous terrain, generally accessible only by foot, intended to provide food and shelter to mountaineers, hikers, and climbers [1]. They are known by many names, such as alpine hut, mountain shelter, mountain refuge, mountain lodge, mountain hostel, and mountain cabin. Mountain huts can provide some functionalities and services. The most important function is to provide shelter and simple sleeping berths for hikers to overcome the cold mountain environment at night. Many mountain huts are not staffed, particularly in remote areas. However, others have staff who provide hot meals, drinks, and other services, such as activity spaces (for holding lectures, etc.), clothing sales, and small items.
Energy supply in high mountains remains an open issue to be solved. Generally, grid connection is impractical, because the establishment and maintenance costs are too high. It also causes many environmental problems, e.g., the destruction of the natural ecology in the mountain. Diesel generators as energy suppliers have been used in high mountains for a long time despite their air pollution and noise problems. Due to the development of renewable energy (such as solar, wind, and hydropower), the usage of diesel generators is reduced, lowering the emissions of greenhouse gases (GHGs). However, owing to their fluctuating nature, most renewable energy sources exhibit intermittent features. To deal with the problem of renewable energy intermittency, energy storage systems (ESSs) are necessary [2, 3]. In the past, lead-acid batteries were heavily used as ESSs, accompanied by many environmental issues, e.g., poisoning, Pb leaks, contamination of the environment, and damage to the ecosystem [4, 5]. Recently, lithium-ion (Li-ion) batteries [6, 7], as a greener alternative, have started to replace lead-acid batteries in ESSs.
Jiujiu Cabins, a famous mountain hut located in Shei-Pa National Park, Taiwan, has adopted an off-grid solar ESS to provide electricity. However, lead-acid batteries were implemented as ESS, accompanied by many environmental issues. In 2021, owing to the aging of lead-acid batteries and equipment failures, electricity was unavailable for a short period. After a detailed on-site survey, a solar power system improvement project was initiated. The repaired energy system came back to provide electricity normally. Moreover, the lead-acid batteries were partly replaced by eco-friendly lithium iron phosphate (LFP) batteries, forming a hybrid ESS composed of lead-acid batteries and LFP batteries. A better and greener off-grid solar energy storage system has been established.
In this manuscript, a brief introduction of the Jiujiu Cabins is given in Sec. II. In Sec. III, the current status of the energy architecture of the off-grid solar ESS is presented with a detailed description of each component. A historical development of energy systems is listed in Sec. IV. In 2021, Jiujiu Cabins experienced a brief period of power outage, and the investigation results of the system failure are listed in Sec. V. How to overhaul and improve the system, and the engineering work to restore power, is covered in Sec. VI. Possible future system upgrade plans are listed in Sec. VII. Summary and Outlook are given at last (Sec. VIII).
II About Jiujiu Cabins
II-A Current status
Jiujiu Cabins (coordinate: $24^{\circ}28^{\prime}08.7^{\prime\prime}$N, $121^{\circ}12^{\prime}29.9^{\prime\prime}$E) is an accommodation mountain hut in Taiwan, located in Shei-Pa National Park, in Tai’an Township, Miaoli County, at an altitude of 2,699 m (8,855 ft). It is named after the number 99 in its altitude [8]. The cabin complex belongs to the Hsinchu Branch, Forestry and Nature Conservation Agency, Ministry of Agriculture, Taiwan. It includes Chengkung Yurts, a restaurant (reception room), Longmen Inns, a machine room, toilet, etc., all of which are separate buildings (as shown in Fig. 2). Jiujiu Cabins uses a photovoltaic system and is equipped with a lithium-ion battery ESS to cope with the intermittent nature of renewable energy [8]. It can currently accommodate 150 people [9, 10]. It has clean running water (but currently no hot water provided) and a mobile phone communication point [11].
Figure 1: Entrance of Jiujiu Cabins (2021).
Figure 2: Aerial photography of Jiujiu Cabins (2021). The cabin complex. It includes Chengkung Yurts, a restaurant (reception room), Longmen Inns, a machine room, toilet, etc., all of which are separate buildings. (Photographer: Yu-Chun Lin)
II-B Brief history
In the 1960s, Jiujiu Cabins was built with lead sheets in the shape of a yurt, which is a common memory of mountain climbers [12] and was later called “Chengkung Yurt.” These six Chengkung Yurts are made of metal and are round mountain houses in the shape of yurts.
Since 1971, in order to cooperate with the China Youth Corps in organizing the Dabajian Mountain hiking event, the Forestry Bureau (now the Forestry and Nature Conservation Agency, Ministry of Agriculture, Taiwan) has opened a hiking trail about 12 kilometers long from the Dalu Forest Road Madala Creek Trailhead to Dabajian Mountain. In addition, Jiujiu Cabins were built in the mountain valley on the west side of the shoulder ridge in the northwest of Jiali Mountain to provide accommodation for the hiking team [13, 14, 15]. To date, there are 6 Chengkung Yurts and 3 Longmen Inns as accommodation spaces, all of which are dormitory-style (with soft mattresses). Hikers need to bring their own sleeping bags and other equipment [8].
II-C Geography
Jiujiu Cabins are an important place for mountain enthusiasts to stay and rest on their way to Daba Peaks (Dabajian Mountain, Xiaobajian Mountain, Yizhe Mountain, Jiali Mountain, etc.) [16]. They usually drive from Dalu Forest Road to Guanwu National Forest Recreation Area [12], and then hike along the flat forest road to the 19 km mark of Madala Creek Trailhead, and then continue for another 4 km along the mountain trail to reach Jiujiu Cabins [17].
III Architecture of current energy system
Due to the difficulty of power grid connection in high-altitude areas, energy supply remains a challenging problem. Off-grid solar ESSs [18, 19] provide an appropriate solution for using renewable energy in high-altitude areas and have been proven [18].
In order to restore the operation of the energy system, the Hsinchu Forest District Office, Forestry Bureau (renamed as the Forestry and Nature Conservation Agency in 2023) commissioned Super Double Power Technology Co., Ltd. to improve the existing solar energy system of Jiujiu Cabins in 2021 [20]. After restructuring and improvement, the off-grid solar energy storage system of Jiujiu Cabins is composed of two independent power supply systems combined with the same energy storage system architecture, as shown in Fig. 3. Among them, the first PV array inputs solar energy to hybrid solar inverters, which can transmit power to alternating current (AC) loads. The second PV array can transmit power to another set of AC loads through another hybrid solar inverter. The two sets of hybrid solar inverters share the same hybrid ESS, which consists of a newly established LFP battery [21] ESS and an existing lead-acid battery ESS. In addition, a new backup power input terminal is added, which can be connected to an external generator as a backup power source. The details of the energy architecture are described below.
Figure 3: Energy architecture of Jiujiu Cabins (2021). The off-grid solar ESS of Jiujiu Cabins is composed of two independent power supply systems combined with the same hybrid ESS. This ensures that two PV arrays can generate electricity and supply two sets of 110 VAC loads. The ESS is a newly added LFP battery system, which is used in conjunction with the existing lead-acid battery ESS to form a hybrid ESS. An additional backup power input terminal is also added, allowing for the connection of an external generator as a backup power source.
III-A Photovoltaic systems (PV systems)
A photovoltaic (PV) system, as a renewable energy system, converts sunlight directly into direct current (DC) electricity using solar panels.
In Jiujiu Cabins, there are two PV arrays, providing a total power generation capacity of 5.3 kWp. The first channel of the PV array is a ground-mounted PV system with a power generation capacity of 2.1 kWp. It is composed of 28 solar modules, each with a capacity of 75 Wp (SIEMENS Solar module SP75), as shown in Fig. 4. The second channel of the PV array is a ground-mounted PV system with a power generation capacity of 3.2 kWp. It is composed of 40 solar modules, each with a capacity of 80 Wp (SHARP Solar Module NE-80E2E), as shown in Fig. 5.
Figure 4: Ground-mounted PV system 1. PV array 1 is a ground-mounted PV system with a power generation capacity of 2.1 kWp. It is composed of 28 solar modules, each with a capacity of 75 Wp.
Figure 5: Ground-mounted PV system 2. PV array 2 is a ground-mounted PV system with a power generation capacity of 3.2 kWp. It is composed of 40 solar modules, each with a capacity of 80 Wp.
III-B Hybrid solar inverters
The hybrid solar inverter serves the crucial role of linking PV systems, ESS, AC loads, and backup power sources. The total nominal power of the hybrid solar inverter system is 12 kW.
Some features of the hybrid solar inverters are listed: (1) detachable liquid crystal display (LCD) control Panel, (2) built-in Wi-Fi function for APP monitoring, (3) zero (0 ms) transfer time, the best protection for servers, (4) pure sine wave inverter with built-in maximum power point tracking (MPPT) solar charger, (5) selectable power charging current, (6) configurable AC/Solar input priority via LCD setting, (7) auto-restart while AC is recovering, (8) overload and short circuit protection, (9) cold start function, and (10) optional parallel operation of up to 9 pcs.
III-C LFP energy storage system (ESS)
ESS serves a role to balance the intermittent renewable energy.
The LFP batteries [21] are used with BMS in the energy storage system. The batteries are in series connection (16S configuration), making a battery system with a nominal voltage of 48 VDC and a capacity of about 10 kWh. A balance circuit is also applied in the system for balancing the voltage of each battery cell. The information of the ESS can be sent to the EMS via the RS-485 communication port.
III-D Battery management system (BMS)
To use the battery safely, a well-designed BMS is required. A BMS is any electronic system that manages a battery cell, pack, or module system, such as protecting the battery from operating outside its safe operating area, monitoring its status, calculating secondary data, reporting that data, controlling its environment, and cell balancing it [19].
III-E Loads
Two channels of AC output with a voltage of 110 VAC and a frequency of 60 Hz are used in the cabins. Each channel is provided by one solar hybrid inverter. Because no communication between the two inverters, their AC output phases are independent.
III-F Backup power system
A channel with a breaker for a backup power system is reserved. Currently, no power system is connected.
A diesel generator could be the backup power system. When the power from the PV system is insufficient (e.g., rainy days or cloudy days), the diesel generator can supply the insufficient power to the cabins. When the system encounters failures (e.g., temporary shutdown of the hybrid solar inverter), the diesel generator can supply the full power to the cabins.
III-G Energy management system (EMS)
A basic functionality of an energy management system (EMS) is to monitor and record the energy input and output.
Due to limited funds, the total voltage, current input/output, and power input/output of the ESS are monitored. The energy information, such as energy generation, consumption, and cell voltages, is not available. It’s difficult to realize the energy input and output of the system quantitatively.
On the other hand, the network communication is not stable. The signal was intermittent. However, it can still be a reference for whether a system is functioning.
Figure 6: Machine room of Jiujiu Cabins (2021).
IV Historical development of energy systems
According to data from the Government e-Procurement System [22], although the Hsinchu Forest District Office commissioned Sheng Yuan Construction Co., Ltd. to carry out a solar panel improvement project in 2003 [23], the system may have undergone several changes over the past decade, and it is impossible to know which companies handled it. This has led to the system becoming complex and difficult to maintain.
The Jiujiu Cabins had two existing electrical systems, designated as Old System 1 and Old System 2. These systems were likely installed by different manufacturers at different times. Owing to a malfunction in the Old System 1, power was unavailable. Another manufacturer then installed Old System 2 to supply power to the cabins’ 110 VAC loads. However, the inverter in Old System 2 also failed due to aging, causing power supply anomalies and no outputting electricity.
In order to restore the operation of the energy system, the Hsinchu Forest District Office commissioned Super Double Power Technology Co., Ltd. to reorganize, repair, and improve the existing solar energy system of Jiujiu Cabins in 2021 [20]. After the improvement project, it’s the current energy architecture (which is described in Sec. III in detail).
The evolution of the energy system in Jiujiu Cabins is shown in Table I.
Year Renewable energy Energy storage system Load capacity Energy management system Project
2021 Total: 5.3 kWp (PV) Total: 25 kWh (Hybrid) 12 kW Yes [20]
Component: Component:
Ch 1: 2.1 kWp (PV) New LFP: 10 kWh
Ch 2: 3.2 kWp (PV) Old lead-acid: 15 kWh
Before Total: 5.3 kWp (PV) 15 kWh (Lead-acid) 3 kW No
2021
Component:
Ch 1: 2.1 kWp (PV)
Ch 2: 3.2 kWp (PV)
TABLE I: Evolution of the Energy System of Jiujiu Cabins
V On-site survey of energy systems
The aging of the energy system caused multiple system failures. In order to understand the energy system status of Jiujiu Cabins, from May 12 to 13, 2021, the Hsinchu Forest District Office, “Wei-Chih, CHEN Architect & Associates,” and Super Double Power Technology Co., Ltd. conducted an on-site survey at Jiujiu Cabins.
The Jiujiu Cabins has two power systems, namely Old System 1 and Old System 2, as shown in Fig. 7. These two systems were likely installed by different manufacturers at different times.
Figure 7: Old energy architecture of Jiujiu Cabins (2021-05-13). The solar charger, lead-acid battery, and inverter of the Old System 1 all failed. Therefore, the Old System 2 was used instead, but the inverter of the Old System 2 also failed, resulting in a complete power supply abnormality and inability to output power.
The Old System 1 can be considered as a microgrid system built on a 48 VDC DC bus. Solar panels (2.1 kWp) generate electricity, which is then charged via a solar charger (PWM) to a 48 VDC lead-acid battery ESS. A 48 VDC inverter converts the DC power to 110 VAC for output. Based on the property label, its installation date is Jan. 20, 2000.
The Old System 2 can be considered as a microgrid system built on a 24 VDC DC bus. After generating electricity from the solar panels (3.2 kWp), the electricity is charged to the 24 VDC lead-acid battery ESS through a solar charger (PWM). The DC power is then converted to 110 VAC AC power output by a 24 VDC inverter, with a rated power of 3 kW.
Fault Status of Old System 1:
  1. 1.
[…]
  2. 2.
    48 VDC inverter failure: Unable to output 110 VAC power.
  3. 3.
    Solar charger failure.
Fault Status of Old System 2:
  1. 1.
    The 24 VDC inverter is faulty and unable to output 110 VAC power.
  2. 2.
    When the No Fuse Breaker (NFB) state is OFF, the solar charger and inverter are connected via a circuit, potentially causing mutual interference. This is due to poor system circuit design.
VI Solar power system improvement project (2021)
In 2021, the Hsinchu Forest District Office commissioned Super Double Power Technology Co., Ltd. to carry out a solar energy system improvement project [20] to solve the system failure issues. The project took place in August 2021 (starting in August and completed in August).
VI-A Key issues
The following are the key issues that need to be addressed:
  1. 1.
    Interference may occur between the two old systems under certain operating conditions, which is one of the reasons why the systems are prone to failure later. The two old systems need to be reassembled into one system.
  2. 2.
    The original inverters are all damaged and need to be replaced.
  3. 3.
    The two existing PV arrays can be reused, but the series and parallel connections need to be adjusted within the inverter specifications.
  4. 4.
    The original lead-acid battery ESS is aging and has insufficient capacity, requiring an increase in the ESS’s capacity to ensure nighttime power supply.
  5. 5.
    The power lines need to be rebuilt.
VI-B Main engineering subprojects
The solution involves restructuring the solar power lines, upgrading the inverters, restructuring the power supply lines at the cabins, and increasing the capacity of the ESS. However, this is not a simple matter of adding more batteries in parallel. It involves the parallel connection of old and new ESSs and the integration of the entire energy system.
The main engineering subprojects are: (1) adding LFP battery modules, (2) updating inverters, (3) modifying and reorganizing the existing PV array 1, (4) modifying and reorganizing the existing PV array 2, (5) changing control boxes (changing AC/DC lines), and (6) materials (hardware parts/connectors/DC lines). The following describes each subproject in detail:
  1. 1.
    Adding LFP battery modules. The original lead-acid battery pack was 24 VDC / 600 Ah. After being modified by connecting it in series and parallel, it was reassembled into a 48 VDC / 300 Ah battery pack with a total capacity of approximately 15 kWh. Two new 48 VDC / 100 Ah LFP batteries were added, with a total capacity of approximately 10 kWh. The modified lead-acid battery pack (15 kWh) and the new LFP battery pack (10 kWh) were connected in parallel, for a total capacity of approximately 25 kWh (neglecting the aging of the lead-acid battery).
  2. 2.
    Updating inverters. The damaged inverter (24 VDC / 3000 W) was removed. Two new independent hybrid solar inverters (48 VDC / 6000 W) were installed. After installation, the total output AC power is 12 kW.
  3. 3.
[…]
  4. 4.
[…]
  5. 5.
    Changing control boxes (changing AC/DC lines). (a) Two damaged solar chargers were removed; (b) the damaged inverter was removed; (c) the AC/DC circuit was modified; and (d) a generator circuit was added.
  6. 6.
    Materials (hardware parts, connectors, and DC lines). (a) The old hardware parts, connectors, and DC lines were removed; (b) new hardware parts, connectors, DC lines were added according to the current decoration situation, including DC surge absorbers, DC dedicated leakage circuit breakers, coulomb meters, cloud EMS, and generator charging input terminals.
VI-C EMS data analysis
In this project, due to budget constraints, the self-developed cloud-native EMS only records the total voltage, input/output current, and input/output power of the ESS. Because the cabins’ energy usage has a certain cyclical characteristic on a daily basis, a one-day data analysis is used to assess the system’s condition, as shown in Fig. 8.
[…]
The 24-hour data shows that the highest total voltage of the ESS does not exceed 54 VDC (as shown in Fig. 8(a)), indicating that the BMS provides stable control over the LFP battery, keeping it within a safe operating voltage range (the rated voltage of this ESS is 48 VDC).
Around 6:00 AM, a while after sunrise, the PV systems charge the ESS, and the current display shows a positive value as shown in Fig. 8(b), indicating charging (a negative current indicates discharging). As the sunlight gradually intensifies, the charging current increases, reaching a maximum of 46 ADC around 9:15 AM. However, after 9:40 AM, because the ESS is fully charged, the system stops charging, and the current can be seen to drop instantly to 0 ADC. At this point, the cabins’ electricity is entirely supplied by solar energy.
Around 5 PM, solar power generation weakened to the point where it was insufficient to supply the cabins’ electricity. The shortfall was covered by the ESS, and the discharge current of the ESS could be seen gradually increasing. By around 7 PM, solar power had completely stopped supplying electricity, and the ESS was fully engaged to power the cabins, resulting in a noticeable drop in the discharge current. The cabins’ electricity consumption peaked around 8 PM, at approximately 620 W (as shown in Fig. 8(c)), and continued until nearly 6 AM the next morning when solar power began generating electricity to support part of the cabins’ needs, at which point the discharge current of the energy storage system gradually decreased. The above is a 24-hour electricity consumption cycle record and analysis of the Jiujiu Cabins, demonstrating the stable operation of the system.
VII Future plan for the energy system
After completing the aforementioned improvement project in 2021, architect Wei-Chih Chen asked Super Double Power Technology to design the future energy system of Jiujiu Cabins.
This plan can be divided into four main components: (1) replacing the old PV arrays with a 20 kWp PV system, (2) replacing the environmentally unfriendly lead-acid battery ESS (15 kWh) with the original LFP ESS (10 kWh), and adding an additional 80 kWh of LFP ESS, (3) installing a self-developed cloud-native EMS, and (4) integrating AC and DC systems. They are stated in detail below.
VII-A PV system replacement
Key engineering points are listed:
  1. 1.
    The existing PV modules of low efficiency (two channels, totaling 5.3 kWp) should be dismantled and removed. Existing usable brackets and fasteners should be collected and retained.
  2. 2.
    New PV arrays should have a capacity of 20 kWp or higher, and the PV modules should be Class-A with a module efficiency of 20% or higher. Product verification conforming to IEC performance and safety standards should be obtained.
  3. 3.
    Solar DC power protection devices such as surge absorbers, lightning bypass, short-circuit protectors, and overload circuit breakers should be installed in the solar DC power control enclosure.
  4. 4.
    Existing aluminum PV brackets should be assessed for any corrosion, paint peeling, cracking, deformation, or other conditions affecting structural strength, and a design should be implemented to maximize their retention and reuse to achieve environmental protection goals.
  5. 5.
    Outdoor connections to PV power lines should have sufficient high-voltage insulation and waterproofing capabilities. For example, the Stäubli MC4 solar-specific waterproof connector can be used, which has 1000 V high voltage and IP68 waterproofing capabilities.
  6. 6.
    To withstand the extremely harsh climate of high-altitude, low-temperature environments, the silicone material used for filling gaps should possess excellent weather resistance, maintaining its elasticity for extended periods. It should be ozone-resistant, ultra-violet-resistant (UV-resistant), and free from aging, cracking, or damage. It should be moisture-resistant and corrosion-resistant, preventing severe electrolytic corrosion. It should withstand temperature variations between freezing and boiling points while retaining the properties of the raw materials. It should have good adhesion and remain bonded for a long time without detaching.
  7. 7.
    During construction, a diesel generator paired with an ESS can be used to supply electricity to the cabins, ensuring a normal power supply of the energy system.
VII-B ESS replacement and enlargement
Key engineering points are listed:
  1. 1.
    The existing ESS comprises two groups: one is a 15 kWh lead-acid battery ESS, which is to be phased out due to battery aging and poor energy storage performance; the other is a 10 kWh LFP ESS built in 2021, which will be retained and integrated with the newly purchased energy storage equipment in this plan.
  2. 2.
    The newly constructed ESS uses an 80 kWh LFP ESS with a rated voltage of 48 VDC and integrates with the operation of the existing ESS.
  3. 3.
    The ESS cabinet can adopt a modular design to distribute the cabinet weight. This allows the mountain porters to successfully bring the materials up the mountain.
  4. 4.
    The battery can adopt a modular design to distribute the battery weight. This allows the mountain porters to successfully bring the materials up the mountain.
VII-C EMS development
The new energy system is expected to have approximately 40 energy nodes. An embedded system will collect, calculate, statistically analyze, and store data from these nodes. All energy data will be uploaded to a cloud-based energy management system via an industrial-grade 4G LTE router.
This plan will integrate hardware and software while preserving the legacy system architecture to the greatest extent possible. The main tasks of the cloud-based EMS are:
  1. 1.
    Hardware integration of the ground system
  2. 2.
    Software integration of the ground system
  3. 3.
    Software integration of the cloud system
The three main work projects are explained as follows.
VII-C1 Hardware integration of the ground system
The points to note are listed below:
  • Install 7 smart DC meters (solar + DC load). Meter accuracy should be Class-0.5 or higher.
  • Install 3 smart AC meters (220 VAC output). Meter accuracy should be Class-0.5 or higher.
  • Install 1 smart AC meter (110 VAC output). Meter accuracy should be Class-0.5 or higher.
[…]
  • Install 4 industrial-grade embedded systems (for ESS). Requirements are as stated in the previous point.
  • Add a new router.
  • Reorganize network, RS-485, RS-232, and CAN Bus communication lines.
  • Reorganize the power lines for communication devices.
VII-C2 Software integration of the ground system
The points to note are listed below:
  • For the 40 newly added energy nodes, integrate ground-level energy information, including voltage, current, power, etc. (information listed below).
  • Decode the new protocol of ESS and calculate energy data, and develop embedded system software for computation.
The energy information from the input and output terminals is listed below:
  • PV input data: voltage, current, power, and total energy generation.
  • 220 VAC input/output data: voltage, current, power, power factor (PF), total energy generation/consumption.
  • 110 VAC output data: voltage, current, power, PF, total energy consumption.
Energy statistics items are listed below:
  • Daily total PV energy generation, daily total energy consumption of the cabins, and daily energy generation of the generator.
  • Monthly total PV energy generation, monthly total energy consumption of the cabins, and monthly energy generation of the generator.
  • Annual total PV energy generation, annual total energy consumption of the cabins, and annual energy generation of the generator.
VII-C3 Software integration of the cloud system
In terms of software and system design, there are several points to note:
  • Integrate cloud-native energy information for over 40 energy nodes.
  • Design an energy information dashboard for the management station. Design an energy information dashboard for engineers.
  • Increase database space to handle the increased data volume.
  • Increase the number of cloud server snapshots, maintaining at least two snapshots to ensure data maintenance.
  • Add a second set of cloud servers and databases to maintain the stability of cloud data delivery.
Regarding cloud environments, there are several points to note:
  • Computing units should be hosted on a reliable and scalable infrastructure. A Service-Level Agreement (SLA) commitment with 99.99% availability is required, with capacity scalable within minutes. A secure computing environment should also be provided.
  • High-performance block storage is an easy-to-use, scalable, designed for use with computing units. It features rapid scaling to meet the most demanding high-performance workloads, including mission-critical applications such as SAP, Oracle, and Microsoft products. It prevents failures with high availability, including write-through within available zones (AZs) and 99.999% durability using io2 Block Express disks. The disks offer high performance with high IOPS and throughput.
  • Virtual cloud environments allow administrators to define and launch cloud resources within logically isolated virtual networks. Administrators have complete control over the virtual network environment, including resource placement, connectivity, and security. This is achieved by configuring the virtual cloud environment in the service console. Next, computing units and high-performance block storage are configured. Finally, the communication between virtual cloud environments, across accounts, and within availability zones is defined for zone communication.
  • The load balancer protects applications using SSL/TLS termination, integrated credential management, and client-side credential verification. It provides highly available and automatically scaling applications. It monitors application operational status and performance in real time, identifies bottlenecks, and maintains SLA compliance.
  • The data storage service used should be scalable, data available, secure, and efficient. It should protect any amount of data, such as data lakes, websites, cloud-native applications, backups, archives, machine learning, and analytics. Its storage service should be designed to achieve 99.999999999% (11 nines) durability.
To understand the energy status of the system, it is essential to be able to grasp the energy input and output. For input, please refer to the energy input line monitoring settings as shown in Table II; for output, please refer to the energy output line monitoring settings as shown in Table III.
No. Line name Description Nominal voltage
1 PV1 PV array 1 <600<600 VDC
2 PV2 PV array 2 <600<600 VDC
3 PV3 PV array 3 <600<600 VDC
4 PV4 PV array 4 <600<600 VDC
5 PV5 PV array 5 <600<600 VDC
6 PV6 PV array 6 <600<600 VDC
7 Generator Diesel generator 220 VAC
TABLE II: Energy input line monitoring settings
No. Line name Description Nominal voltage
1 LineGeneral Typical 110 VAC power consumption 110 VAC
2 Line220VAC Dedicated for 220 VAC loads 220 VAC
3 Line48VDC Dedicated for 48 VDC loads 48 VDC
4 LineEMS Monitoring the power consumption of the EMS 220 VAC
TABLE III: Energy output line monitoring settings
VII-D AC and DC systems integration
This project involves both AC and DC systems within a microgrid. The power lines for these two systems should be independent; otherwise, system damage may occur. Appropriate physical isolation (e.g., breakers, switches, etc.) is required between each device to facilitate installation and subsequent maintenance. Other points to note:
  1. 1.
    The PV system, ESS, and EMS mentioned above need to be integrated, with parameters matched to ensure they can operate collaboratively.
  2. 2.
    The existing inverters should be used in conjunction with the newly installed inverters, operating in parallel to supply power to the cabins.
  3. 3.
    The inverter’s DC operating voltage setting should match the operating voltage setting of the LFP ESS to prevent the inverter from draining the battery’s power.
  4. 4.
    When connecting different batteries in parallel, the system voltage difference must be reduced before connecting them. Excessive voltage difference can cause instantaneous currents of hundreds to thousands of amperes, damaging the system.
After system adjustments, an off-grid solar ESS will be established to achieve the purposes of energy generation, storage, and consumption. Solar energy will be the primary energy source, while the diesel generator will serve as a backup. The LFP ESS will be used for energy storage, supplying energy to the cabins’ loads via hybrid solar inverters. The system architecture is shown in Fig. 9. On the left side, at the energy input, PV arrays will be the primary energy source, while the generator will be an auxiliary system, supplementing the required power when the solar energy supply is insufficient or malfunctions. On the right side, at the energy output, power will be provided for 220 VAC loads, 110 VAC loads, and 48 VDC loads.
Junction boxes are divided into three main categories: PV panel, AC panel, and DC panel. The PV panel provides DC power transmission between the PV arrays and hybrid solar inverters. The AC panel provides power transmission from the hybrid solar inverters to 220 VAC loads, from the generator to the hybrid solar inverters, and from the 220 VAC power to the AC transformer, converting the 220 VAC power to supply 110 VAC loads. The DC panel provides bidirectional power transmission between the hybrid solar inverters and the LFP ESS.
Additionally, a self-developed cloud-native EMS can detect PV power generation (unit: W), PV energy generation (unit: kWh), load AC power consumption (unit: W), and load AC energy consumption (unit: kWh), showing the values on the enclosure using meters with liquid crystal display (LCD). The information is also uploaded to a cloud server for remote monitoring.
Figure 9: Future energy architecture of Jiujiu Cabins (planned in 2021). On the energy input side, PV arrays will be the primary energy source, while the generator will be an auxiliary system, supplementing the required power when solar energy supply is insufficient or malfunctions. On the energy output side, power will be provided for 220 VAC loads, 110 VAC loads, and 48 VDC loads.
VIII Summary and Outlook
Using renewable energy and eco-friendly ESS in high mountains is not easy. It took about two decades to approach the goal in Jiujiu Cabins. Since 2003, renewable energy provided by PV systems has been used. However, the lead-acid battery ESS still causes pollution in high mountains. Until 2021, encountering a system failure and repairing period, the eco-friendly LFP ESS replaces part of the lead-acid battery ESS.
In the next step, the remaining lead-acid batteries should be replaced by the eco-friendly LFP ESS entirely, making a better and greener off-grid solar ESS. On the other hand, a more comprehensive EMS should be installed to check the energy input and output. Figuring out the energy usage of the cabins. Further suggestions can be given after the energy usage analysis, such as “Is the energy generated sufficient? Is it necessary to expand the PV system?” or “Is the energy stored sufficient? Is it necessary to expand the ESS?”
We wish not only for a green building in high mountains, but also for using green energy to protect the ecosystem and environment.
Acknowledgment
The author (H.-C. Chung) would like to thank the contributors to this article for their valuable discussions and recommendations: Jung-Feng Jack Lin, Hsiao-Wen Yang, Yen-Kai Lo, and An-De Andrew Chung. The author (H.-C. Chung) thanks Pei-Ju Chien for English discussions and corrections, as well as Ming-Hui Chung, Su-Ming Chen, Lien-Kuei Chien, and Mi-Lee Kao for financial support. This work was supported in part by Super Double Power Technology Co., Ltd., Taiwan, under the project “Development of Cloud-native Energy Management Systems for Medium-scale Energy Storage Systems (https://osf.io/7fr9z/)” (Grant number: SDP-RD-PROJ-001-2020).